Operations Report – January 31st

SOL: 10
Name of person filing report: Scott Dorrington
Non-nominal systems: None
Notes on non-nominal systems: None
Spirit rover used: Not used
Opportunity rover used: Not used
Curiosity rover used: Not used
Perseverance rover used: Not used
General notes on rovers:
Summary of Hab operations:
WATER USE: 21.15 gallons
Water (static tank): 223.25 gallons (depth 25”)
Static tank pipe heater (on or off): On
Static tank heater (on or off): On
Toilet tank emptied: no
Summary of internet: Nominal during comms window
Summary of suits and radios:
Batteries on suits #6 and #7 were replaced with new ones. Old batteries were labelled and placed with other faulty batteries. Charged batteries for 1 hour after replacement, then ran fans for 1 hour, testing voltage before and after. The voltage on suit 7 is also now stable after discharging.
Suit 6: Start: 13.60 V, End: 10.78 V (after 1 hr running)
Suit 7: Start: 13.08 V, End: 12.38 V (after 1 hr running)
Suit 8: Used nut kit in the RAM to judge the size of the nut. Found it is bigger than M3 and smaller than M4. Likely an imperial measurement.
Suit 11: While inspecting other suits for charging port nuts, I pulled suit 11 off the bottom rack. One of the screws from the back fell to the floor (likely not secured). Opening the back, I discovered the battery cables were not attached to the battery terminals. Not sure if this is deliberate. Left as found. Replaced screws in back, half screwed in – not all the way tightened.
Summary of GreenHab operations:
WATER USE: 9 gallons
Heater: On
Supplemental light: 10pm – 2am
Harvest: None
Summary of ScienceDome operations: Andrew used Science Dome for processing of gypsum discovered during yesterday’s EVA to White Moon. Gypsum was crushed in mortal and pestle and heated on hotplate to extract water. Clare used Science Dome to wash rock samples from previous EVAs. Steve used Science Dome to gather spectral curves of rock samples. Scott observed activities in Science Dome.
Dual split: Nominal
Summary of RAM operations: Entered RAM on several occasions to retrieve and return tools. Used tin snips to cut up a can for gypsum processing. Used socket sets and nut kit to investigate size of nut used for charging port on one-piece EVA suits.
Summary of any observatory issues: See astronomy report
Summary of health and safety issues:
Crew member received graze on head from hitting it on the door frame of the observatory. Bandages applied. Monitoring.
Questions, concerns and requests to Mission Support:
Advise on state of suit #11. Should the battery terminals be attached and screws inserted.

Copyright © The Mars Society. All rights reserved. | Main Site